OpenStudy (anonymous):

Help ! This makes no sense... Solve by elimination x + 5y - 4z = -10 2x - y + 5z = -9 2x - 10y - 5z = 0 ....

OpenStudy (anonymous):

First, pick two equations to start with

OpenStudy (anonymous):

The first two appears easier to me. Multiply the first one by -2 in order to eliminate the X

OpenStudy (anonymous):

-2x -10y +8z = -20 2x - y + 5z = -9 -9y + 13z = -29 (combine the two equations) Now, do equation 2 and 3 : Multiply original equation 3 by -1 to eliminate the x

OpenStudy (anonymous):

The line abve (combination) should be -11y + 13z = -29

OpenStudy (anonymous):

equation 2 multiplied by -1 -2x + y -5z = 9 2x - 10y -5z = 0 (orig #3) -9y -10z = 9 combined 2 and 3

OpenStudy (anonymous):

A. combined 1 & 2: -11y + 13z = -29 B. combined 2&3: -9y - 10z = 9 multiply A by 9 multiply B by -11 -99y + 117 z = -261 99y + 110z = -99 227z = -360 z = -360/227

OpenStudy (anonymous):

Did you follow? Three equations are long...but has the same logic as 2 equations -

OpenStudy (anonymous):

I'm trying to follow.. Very difficult though.

OpenStudy (anonymous):

I know :(

OpenStudy (anonymous):

labeled the original equations as A B C

OpenStudy (anonymous):

It just doesn't make sense to me. lol

OpenStudy (anonymous):

We are going to canceled out terms with equation A and B Multiply A by -2 Combine the A' with B

OpenStudy (anonymous):

when you combine A' and B you get -11y + 13z = -29

OpenStudy (anonymous):

now Multp B by -1 Combine B' with C you get -9y -10z = 9

OpenStudy (anonymous):

combine those two and solve for Z we get -360/227 I hope my calculations were right..)

OpenStudy (anonymous):

I almost understand. I guess I just need to read those stupid books more.

OpenStudy (anonymous):

Now we start with -9y -10z = 9 plug in z = -360/227 and solve for Y

OpenStudy (anonymous):

:).. let make the fraction of -360/227 a decimal = -1.59

OpenStudy (anonymous):

I used the -9y -10z = 9 to plug for Z and solve for Y I got Y = .76

OpenStudy (anonymous):

now I look at original A plug in my values for Y and Z to solve for X

OpenStudy (anonymous):

Oh dear.. Ohkay.. Let me try lol No promises

OpenStudy (anonymous):

I get x = -20.2

OpenStudy (anonymous):

I can't figure it out. Plus, my calculator won't do decimals. ):

OpenStudy (anonymous):

I am off on the Z I can't get it to check. I am still looking at it. Your calc won't decimals? strange...

OpenStudy (anonymous):

It doesn't have the decimal point on it. It's a mini calculator

OpenStudy (anonymous):

check out this video. same steps as I but maybe fractions work better for you http://www.mathplanet.com/education/algebra-2/how-to-solve-system-of-linear-equations/solving-systems-of-equations-in-three-variables

OpenStudy (anonymous):

Alrighty(:

OpenStudy (anonymous):

Looks a little more simple. Thank you(:

OpenStudy (doc.brown):

\[\left[\begin{matrix}1 & 5 & -4 & -10 \\ 2 & -1 & 5 & -9 \\ 2 & -10 & -5 & 0 \end{matrix}\right]\]\[row_{2} += (-2)row_{1}\]\[\left[\begin{matrix}1 & 5 & -4 & -10 \\ 0 & -11 & 13 & 11 \\ 2 & -10 & -5 & 0 \end{matrix}\right]\]\[row_{3} += (-2)row_{1}\]\[\left[\begin{matrix}1 & 5 & -4 & -10 \\ 0 & -11 & 13 & 11 \\ 0 & -20 & 3 & 20 \end{matrix}\right]\]\[row_{3} += (\frac{-20}{11})row_{2}, += \frac{row_{3}}{a_{33}}\]\[\left[\begin{matrix}1 & 5 & -4 & -10 \\ 0 & -11 & 13 & 11 \\ 0 & 0 & 1 & 0 \end{matrix}\right]\]\[row_{2} += (-13)row_{3}, += (\frac{-1}{11})row_{2}\]\[\left[\begin{matrix}1 & 5 & -4 & -10 \\ 0 & 1 & 0 & -1 \\ 0 & 0 & 1 & 0 \end{matrix}\right]\]\[row_{1} += (-5)row_{2}, += (4)row_{3}\]\[\left[\begin{matrix}1 & 0 & 0 & -5 \\ 0 & 1 & 0 & -1 \\ 0 & 0 & 1 & 0 \end{matrix}\right]\]\[x= -5\]\[y= -1\]\[z= 0\]

OpenStudy (anonymous):

THis is where i erred when you combine A' and B you get -11y + 13z = -29 that should be -11y + 13z = 11

OpenStudy (anonymous):

Wow... That's a lot. Thanks whoever you are. lol

OpenStudy (anonymous):

when you combine -11y + 13z = 11 with 9 y + 10z = -9 by multiply 1st by 9 and the 2nd by 11 you get z = 0

OpenStudy (anonymous):

from that...plug z = 0 into 9y +10z = -9 you get y = -1

OpenStudy (anonymous):

That makes more sense. No decimals.. Thank goodness

OpenStudy (anonymous):

plug y = -1 and z = 0 into original 1st line you get x = -5 yes matrix is a way to do this also...

OpenStudy (anonymous):

I figured I did math error somewhere, but i found it when I went back through my steps. sorry..

OpenStudy (anonymous):

It's fine (: that website you sent helped a lot